# Casimir-Lifshitz interaction between bodies integrated in a micro/nanoelectromechanical quantum damped oscillator

### v1

@@ Abstract
This paper proposes a microscopic theory of Casimir-like forces induced by the quantum damped oscillator. The application of this approach to bodies integrated in micro/nanoelectromechanical systems reveals the conditions for a lumped description of the circuit-induced Casimir-Lifshitz forces. While the free energy diverges under those conditions, the circuit-induced component of the force has been identified and shown to be finite.
@@
The results obtained can be straightforwardly extended to other problems involving damped oscillators or electrical circuits, where a derivative of free energy $F$ with respect to a parameter other than the interbody distance is of interest. The only requirement is that the dependence of $F$ on this parameter is entirely determined by the dependence of $\Omega$ and $\gamma$ on it. For example, an interaction torque arises when $\Omega$ and/or $\gamma$ depend on the misorientation angle $\theta$ between the interacting bodies. This torque can be derived from (1)–(6) by replacing $\frac{\partial\Omega}{\partial d}$ and $\frac{\partial\gamma}{\partial d}$ with $\frac{\partial\Omega}{\partial\theta}$ and $\frac{\partial\gamma}{\partial\theta}$, respectively.
@@
Osipyan Institute of Solid State Physics RAS, Chernogolovka, Moscow District, 2 Academician Ossipyan Street, 142432 Russia
@@
It follows from (S15) that the circuit-induced interaction force for the RLC circuit in series is
$f^{s}_{RLC}=\frac{T}{2C}\frac{\partial C}{\partial d}+\frac{i\hbar\frac{\partial C}{\partial d}}{2\pi C^{2}R\sqrt{\frac{4L}{CR^{2}}-1}}\left[\psi\left(1+\frac{\hbar R}{4\pi TL}\left(1-i\sqrt{\frac{4L}{CR^{2}}-1}\right)\right)-\right.\\ \left.-\psi\left(1+\frac{\hbar R}{4\pi TL}\left(1+i\sqrt{\frac{4L}{CR^{2}}-1}\right)\right)\right],$ (S20)
@@
It can be verified that the factor preceding the term $\frac{\partial\gamma_{0}}{\partial d}$ in (S39) is positive throughout the domain of applicability of the derived expression $\omega_{D}\gg\Omega,\gamma$. Therefore, in accordance with the result stated in the main text, each of the three contributions to the Casimir force in (S39) describes attraction, when the corresponding parameter $\Omega$, $\gamma_{0}$, or $\omega_{D}$ increases with $d$, and repulsion when it decreases with $d$.

### v2

@@ V Conclusions
This paper develops a theory of Casimir-like forces induced by a quantum damped oscillator. Applied to bodies integrated in micro/nanoelectromechanical systems, this theory identifies the conditions, under which a lumped description of circuit-induced Casimir-Lifshitz forces is valid. Although the free energy diverges under these conditions, the circuit-induced component of the force is shown to be finite. The forces in series and parallel RLC circuits are described in detail.
Two small parameters, one geometric and the other given by the ratio of characteristic frequencies, determine the relative magnitude of circuit-induced fluctuation forces compared to the conventional Casimir-Lifshitz forces. Nevertheless, modern high-precision measurements make detailed studies of these forces feasible. Moreover, at high temperatures a single geometric parameter controls the relative magnitude of the effect, and under certain conditions the circuit-induced forces may constitute a substantial fraction of the total force.
@@
The results obtained here can be straightforwardly extended to other problems involving damped oscillators or electrical circuits, in which derivatives of the free energy $F$ with respect to parameters other than the interbody distance are of interest. The only requirement is that the dependence of $F$ on the parameter be entirely determined by the dependence of $\Omega$ and $\gamma$ on it. For example, an interaction torque arises when $\Omega$ and/or $\gamma$ depend on a misorientation angle $\theta$ between the interacting bodies. This torque can be derived from (1)–(6) by replacing $\tfrac{\partial\Omega}{\partial d}$ and $\tfrac{\partial\gamma}{\partial d}$ with $\tfrac{\partial\Omega}{\partial\theta}$ and $\tfrac{\partial\gamma}{\partial\theta}$, respectively.
@@
Osipyan Institute of Solid State Physics RAS, 2 Academician Osipyan Street, Chernogolovka, Moscow Region, 142432 Russia
@@
It follows from (S15) that the circuit-induced interaction force $f^{s}_{RLC}$ for the RLC circuit in series is
$f^{s}_{RLC}=\frac{T}{2C}\frac{\partial C}{\partial d}+\frac{i\hbar\frac{\partial C}{\partial d}}{2\pi C^{2}R\sqrt{\frac{4L}{CR^{2}}-1}}\left[\psi\left(1+\frac{\hbar R}{4\pi TL}\left(1-i\sqrt{\frac{4L}{CR^{2}}-1}\right)\right)-\right.\\ \left.-\psi\left(1+\frac{\hbar R}{4\pi TL}\left(1+i\sqrt{\frac{4L}{CR^{2}}-1}\right)\right)\right].$ (S20)
@@
It can be verified that the factor preceding the term $\frac{\partial\gamma_{0}}{\partial d}$ in (S39) is positive throughout the domain of applicability of the derived expression $\omega_{D}\gg\Omega,\gamma$. Therefore, in accordance with the result stated in the main text, each of the three contributions to the Casimir force in (S39) represents attraction, when the corresponding parameter $\Omega$, $\gamma_{0}$, or $\omega_{D}$ increases with $d$, and repulsion when it decreases with $d$.